OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 13 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

this is easy! eliminate the ys x - y = 13 2x + y = 2 x=13 + 2x=2 3x=15 divide. 3x/=15/3 x= 5 plug it in. (5)-y=13 -5 -5 -y=8 so...x=5 and y=(-8)

OpenStudy (anonymous):

elimination is the mest method because the y is already ready to be crossed out